ON THE DIPOLE MOMENTS OF SOME EXCITED ELECTRONIC STATES OF HfS AND HfO \underline{ANNIE HANSSON}, A. PETTERSSON, P. ROYEN and U. SASSENBERG, Stockholm University, Department of Physics, Stockholm, Sweden.